# Spatial Policy: Guiding Visuomotor Robotic Manipulation with Spatial-Aware Modeling and Reasoning

### v1

@@ Introduction
During visual planning, we incorporate a vision-language model (VLM)-based consistency checker inspired by VideoAgent (Soni et al. 2025). A pretrained GPT-4o model is prompted to assess the spatial plausibility of the generated video—verifying object-goal alignment, manipulator trajectories, and overall spatial coherence. Only videos that pass this check are used for downstream control; otherwise, regeneration is triggered under the same spatial condition.
@@
AVDC 30.7% 28.0% 21.3% 8.0% 34.7% 17.3%
AVDC-R 72.0% 89.3% 37.3% 18.7% 60.0% 24.0%
VA 40.0% 29.3% 13.3% 9.3% 38.7% 18.7%
VA-OR 82.7% 97.3% 40.0% 26.7% 73.3% 44.0%
SP 98.6% 100.0% 66.7% 80.0% 94.7% 96.0%
@@
Method faucet-close faucet-open handle-press hammer assembly Overall
AVDC 12.0% 17.3% 41.3% 0.0% 5.3% 19.6%
AVDC-R 53.3% 24.0% 81.3% 8.0% 6.7% 43.1%
VA 46.7% 12.0% 36.0% 0.0% 1.3% 22.3%
VA-OR 74.7% 46.7% 86.7% 8.0% 10.7% 53.7%
SP 93.3% 89.3% 100.0% 74.7% 24.0% 83.4%
SP-R 92.0% 96.0% 100.0% 73.3% 32.0% 86.7%
Datasets and Experimental Setups
Datasets and Environments. We follow the evaluation setup of (Ko et al. 2023) and conduct experiments in the Meta-World (Yu et al. 2020). Meta-World is a simulated benchmark designed for assessing robotic manipulation capabilities, consisting of 11 representative single-arm tasks (such as pushing, pulling, grasping, and inserting) performed by a Sawyer robotic arm. We run 25 random seeds across 11 tasks and 3 camera views, yielding over 800 rollouts in total, and evaluate 5 ablation studies in our model.
@@
  • AVDC: A two-stage control planning framework is employed, wherein a pretrained diffusion-based generative model synthesizes a video sequence, followed by the estimation of inter-frame motion using an optical flow network. The resulting motion is then used to inversely infer the corresponding action sequence.
  • AVDC-Replan: An extension of AVDC that incorporates a replanning mechanism. During execution, if action failure or robotic arm stagnation is detected, a new video is regenerated and a revised action plan is inferred.
  • VideoAgent: This method employs a self-supervised consistency mechanism to iteratively refine the generated video. Multiple candidate video plans are scored using GPT-4, and the most promising one is selected for execution.
  • VideoAgent-online-Replan: An extension of VideoAgent that continuously collects successful execution trajectories during deployment. These are used to update both the video generation and optimization models online. A replanning strategy is also introduced to regenerate videos when execution failures occur.
  • SP (Ours): We extract structured planning information from spatial context using VLM, and inject it into the diffusion model via global condition to effectively guide video generation. The generated video is subsequently translated into executable actions using a goal-conditioned policy.
  • SP-Replan (Ours): An extended version of our method that performs iterative video generation in the planning phase. During execution, if robotic arm stagnation is detected, a new video is regenerated and the action sequence is re-inferred.
Main Results
As shown in Table 1, our method achieves a success rate of 83.4% without requiring execution-time replanning, already outperforming all prior baselines. In contrast, methods such as AVDC and VA rely on task text to generate videos or pretrained optical flow to compute displacement coordinates for action prediction. These baselines are fragile during deployment, once a deviation or unexpected event occurs, they must fall back on explicit replanning to recover; otherwise, the robot stalls. By contrast, our approach leverages spatial plans to generate structured, goal-aligned videos and predicts executable actions that inherently adapt to spatial variations, leading to more robust and consistent task completion. SP-R further boosts performance to 86.7%, demonstrating the effectiveness of incorporating real-time spatial reasoning into the control loop.
On the subset of challenging tasks ( Assembly, Hammer, Faucet Open, Button Press Top, Shelf Place, and Basketball ), our method achieves an average success rate of 77.5%, significantly outperforming the best-performing baseline, which only reaches 29.2% . This corresponds to a relative improvement of 165.4%. On simpler tasks, our approach achieves near-perfect performance, often approaching 100% success.
Figure 2: Comparative visualization of spatial planning and video generation results for two representative tasks: Shelf Place and Basketball. The yellow line depicts the motion trajectory; the symbol marks the initial position, the symbol indicates the final task position, and the symbols denote intermediate steps along the path.
In the Shelf Place task, as shown in Figure 2, the baseline method generates an incorrect and inflexible trajectory (highlighted in the yellow box), resulting in task failure due to its inability to adapt to environmental variations. In contrast, our approach, enhanced with spatial understanding, enables the robotic arm to follow a dynamically planned path and successfully complete the placement task. Similarly, in the Basketball task (also illustrated in Figure 2), our method exhibits refined spatial reasoning by allocating more frames (highlighted in the green box) around the critical manipulation event—grasping the ball—leading to smoother and more precise motions. The baseline method, lacking this focused spatial adaptation, often misses essential interaction moments. These results underscore the benefits of integrating spatial awareness with dynamic feedback for robust and effective long-horizon robotic planning.
Figure 3: Visualization of the replanning-based spatial planning process for a manipulation task.
Figure 3 further demonstrates how our method leverages spatial feedback and iterative replanning to recover from failures and achieve successful task execution. Unlike baseline planning, which can fail when the initial path is invalid, our approach dynamically adapts, resulting in more robust and reliable performance.
Ablation Study
Ablations on the Video Generation Model. We compare three strategies in the video generation module to evaluate their impact on video quality and downstream control performance. (1) The baseline uses AVDC’s video generation approach, where a diffusion model is conditioned only on task text and the initial frame. (2) The Distance variant directly encodes the spatial offset between the robot and target object (spatial coordinates) and concatenates it with the task prompt as input. (3) The Subplan variant leverages a vision-language model to transform spatial coordinates into structured subplans that encode action type, direction, and distance, which are then injected into the diffusion model.
@@
As shown in Table 2, both spatial variants outperform the AVDC baseline, demonstrating the utility of explicitly modeling spatial structures for enhanced visuomotor performance. Moreover, the subplan-based approach achieves further gains by abstracting raw spatial cues into structured guidance, highlighting the importance of structured spatial abstraction in facilitating precise and actionable visual plan generation.
Ablations on the Action Prediction Model. AVDC uses a pre-trained flow model to estimate pixel-wise motion between the current and goal frames, then converts the flow into a spatial target coordinate. The robot executes actions using a classical controller to move toward the point and applies gripper control based on heuristics. In contrast, our method is built on diffusion policy and directly predicts executable actions from the current frame, goal frame, and robot spatial coordinates. This design captures spatial and temporal context jointly and allows flexible adaptation to unexpected events such as object shifts or failures. Results in Table 3 show that our approach achieves higher success rates and better robustness, confirming that learning-based spatial reasoning provides more effective action prediction than coordinate-based control.
Ablation on Spatial Condition Injection Strategy in Video Generation. To investigate how spatial conditions affect video generation, we conduct an ablation study comparing two injection strategies for integrating the spatial plan table into the diffusion model. (1) The spatial plan table is concatenated with the task embedding and propagated uniformly across all layers of the U-Net backbone in the Global Cond variant directly. This approach provides consistent spatial guidance throughout the generation process. (2) Inspired by GLIGEN (Li et al. 2023), the spatial plan is injected only into intermediate U-Net layers via a cross-attention mechanism in the Local Cond variant, allowing localized spatial modulation while preserving the task-level prompt as a global condition.
Both strategies are trained under identical configurations and evaluated on MetaWorld. As shown in Table 4, Local Conditioning slightly improves the average task success rate and shows better performance on harder tasks (e.g., +5.3% on Hard). These results suggest that injecting spatial structure via localized attention can enhance spatial precision during video synthesis, thereby improving policy success rates, especially in more challenging environments.
Easy Medium Hard Very Hard Overall
@@
Subplan 96.8% 63.3% 24.0% 78.7% 82.4%
Table 2: Comparison of task success rates (%) using different video generation strategies across four levels of task difficulty. AVDC uses text and initial frame only; Distance adds raw spatial offsets; Subplan injects structured spatial subplans.
Easy Medium Hard Very Hard Overall
Flow 25.9% 10.7% 5.3% 8.0% 19.6%
@@
Table 3: Comparison of task success rates (%) between our method and AVDC’s action prediction module across four levels of task difficulty. Flow denotes the flow model-based action prediction used in AVDC; Ours denotes our spatial-based action prediction module.
Easy Medium Hard Very Hard Overall
Global Cond 96.8% 63.3% 24.0% 78.7% 82.4%
@@
Table 4: Comparison of task success rates (%) using different spatial condition injection strategies for video generation across four levels of task difficulty. Global Cond injects the spatial plan table across all U-Net blocks; Local Cond follows GLIGEN-style partial injection via cross-attention at intermediate layers.
Ablations on the Replanning Mechanism. We compare three settings to evaluate the effect of the replanning mechanism: (1) no replanning at all, (2) replanning only during the video generation stage, and (3) replanning during both the video generation and real execution stages. As shown in Table 5, task success improves when replanning is introduced.
Our experiments demonstrate that enabling replanning during both the imagination and real execution stages significantly enhances task success by facilitating dynamic adaptation to the evolving environment. While replanning solely during the imagination stage offers limited improvements, the integration of a vision-language-based evaluation module allows the system to selectively regenerate and retain higher-quality trajectories, effectively filtering out suboptimal predictions despite the inherent stochasticity of video generation. Without any replanning, the model’s performance deteriorates due to its inability to recover from unforeseen variations. Therefore, a combined replanning strategy is crucial for achieving robust and effective long-horizon robotic control.
Easy Medium Hard Very Hard Overall
NR 96.8% 63.3% 24.0% 78.7% 82.4%
@@
Table 5: Comparison of task success rates (%) using different replanning mechanisms across four levels of task difficulty. NR denotes no replanning; GR denotes generation replanning; ER denotes execution replanning. Bold indicates best.
Ablations on Finetuning with Additional Execution Data. To assess the impact of trajectory diversity on policy robustness, we compare two training data configurations for our model. (1) The model is trained exclusively on demonstrations where tasks are completed successfully from start to finish without failure. (2) The model is further finetuned on an expanded dataset that includes recovery trajectories where execution initially failed but eventually succeeded when resumed from intermediate states.
Finetuning on these recovery trajectories yields noticeable gains, particularly on challenging tasks. On btn-press-top, the success rate improves from 97.3% to 99.0%. More strikingly, on the more complex hammer task, performance increases from 68.0% to 78.7%. These results indicate that including partial-to-successful executions enhances the model’s robustness by exposing it to diverse recovery scenarios. Such data enables better generalization to states not frequently seen in the original dataset of consistently successful trials.
Hyperparameter Analysis
@@
Effect of Hyperparameter in Action Prediction. We analyze key design choices in the action prediction module. First, we study the effect of goal frame selection range. Setting the goal frame as a random future frame within the next $k=20$ timesteps yields over 10% higher average success rates compared to restricting it to a narrower window between 17–20 timesteps ahead, indicating that allowing a broader temporal horizon improves generalization and task adaptability. Second, regressing the noise $\epsilon$ achieves better performance than directly predicting the denoised action $\hat{a}_{0}$, which contrasts with some prior findings (Chi et al. 2023). Lastly, while direct MLP-based action regression converges faster, the diffusion-based denoising model demonstrates superior performance and robustness across tasks.
Conclusion
We present SP, a spatial-aware visuomotor framework that integrates explicit spatial modeling and reasoning into robot manipulation. Central to our approach is a spatial plan table that encodes atomic actions, directions, and distances, enabling structured spatial guidance throughout the pipeline. Our polciy generates spatially coherent visual trajectories by conditioning a video diffusion model on spatial plan table, , which are then used to guide action prediction based on temporally aligned spatial coordinates. A dual-stage replanning mechanism enhances spatial reasoning by refining the spatial plan table and regenerating video and actions upon inconsistency, ensuring closed-loop correction. Experiments on MetaWorld show that SP improves success rates by over 33% compared to prior methods, validating the effectiveness of spatial modeling and reasoning.
Appendix A Inference Pipeline

### v2

@@ 3.3 Spatial Reasoning Feedback Policy
During visual planning, we incorporate a vision-language model (VLM)-based consistency checker inspired by VideoAgent [videoagent]. A pretrained GPT-4o model is prompted to assess the spatial plausibility of the generated video—verifying object-goal alignment, manipulator trajectories, and overall spatial coherence. Only videos that pass this check are used for downstream control; otherwise, regeneration is triggered under the same spatial condition.
@@ 4 Experiments
AVDC [avdc] 30.7% 28.0% 21.3% 8.0% 34.7% 17.3%
AVDC-R [avdc] 72.0% 89.3% 37.3% 18.7% 60.0% 24.0%
VA [videoagent] 40.0% 29.3% 13.3% 9.3% 38.7% 18.7%
VA-OR [videoagent] 82.7% 97.3% 40.0% 26.7% 73.3% 44.0%
SP 100.0% 100.0% 77.3% 78.7% 96% 98.6%
@@
[…]
Table 2: Comparison of task success rates (%) across 12 tasks in four rooms on iTHOR.
Room AVDC [avdc] VideoAgent [videoagent] SP (Ours)
Kitchen 26.7% 28.3% 88.3%
Living Room 23.3% 26.7% 50.0%
Bedroom 38.3% 41.7% 63.3%
Bathroom 36.7% 40.0% 36.7%
Average 31.3% 34.2% 59.6%
In this section, we comprehensively evaluate our method across both manipulation and navigation tasks. Sec. 4.1 introduces the environments and datasets, and Sec. 4.2 reports the main results on Meta-World and iTHOR. Sec. 4.3 presents the real-world evaluation results. We further conduct five ablation studies in Sec. 4.4, examine failure modes in Sec. 4.5, analyze computational efficiency in Sec. 4.2, assess robustness to occlusions in Sec. 4.6, and study key hyperparameters in Sec. 4.7.
4.1 Datasets and Experimental Setups
Datasets and Environments. We follow the evaluation setup of [avdc] and conduct experiments in both Meta-World [metaworld] and iTHOR [kolve2017ai2thor]. Meta-World is a simulated robotic manipulation benchmark consisting of 11 single-arm control tasks (e.g., pushing, pulling, grasping, inserting); we run 25 random seeds across 11 tasks and 3 camera views, yielding over 800 rollouts. iTHOR is an interactive 3D household environment featuring four room types and 12 navigation tasks, where each task is evaluated with 20 random seeds to assess robustness and cross-scene generalization. Complete training details are provided in Appendix A.2, generated results in Appendix A.3, and dataset details in Appendix A.4.
@@ 4.1 Datasets and Experimental Setups
  • AVDC: Generates videos via a pretrained diffusion model, extracts optical-flow motion, and inversely infers actions.
  • AVDC-Replan: Adds replanning by regenerating videos when failures or stagnation occur.
  • VideoAgent: Refines video plans with a self-supervised consistency rule and selects candidates via GPT-4 scoring.
  • VideoAgent-online-Replan: Online-updates video generation and optimization using successful trajectories and triggers replanning upon failures.
  • SP (Ours): Extracts structured spatial plans with a VLM and injects them as global conditions into the diffusion model; actions are derived via a goal-conditioned policy.
  • SP-Replan (Ours): Adds iterative regeneration and replanning when execution stagnation is detected.
4.2 Main Results
Meta-World
As shown in Table 1, our method achieves an overall success rate of 86.7%, outperforming all prior baselines. Approaches such as AVDC and VA rely on task text or pretrained flow to infer motion, making them brittle under execution drift and forcing explicit replanning when deviations occur. In contrast, our spatial-plan-driven framework produces structured, goal-aligned video predictions and actions that inherently adapt to spatial variations, resulting in more stable and consistent execution. On the subset of challenging tasks (Assembly, Hammer, Faucet Open, Button Press Top, Shelf Place, Basketball), our method reaches 77.5% success, whereas the best baseline achieves only 29.2%, corresponding to a 165.4% relative improvement. Easier tasks are solved nearly perfectly, often approaching 100% success.
Figure 2: Comparisons on Shelf Place and Basketball in Meta-World. The yellow line shows the motion trajectory; marks the start, the final position, and intermediate steps.
Figure 3: Comparisons on Painting and Mirror in iThor. The bounding box marks the target object.
Fig. 2 illustrates these advantages. In Shelf Place, the baseline produces an inflexible trajectory that cannot correct for scene variations, whereas our model generates adaptive, spatially coherent paths. In Basketball, our method allocates more frames to the critical grasping phase (green box), yielding smoother and more reliable manipulation, while baseline methods frequently miss key interaction moments.
Figure 4: Example of iterative spatial replanning during manipulation in Meta-World.
[…]
iTHOR
Table 2 shows that our method reaches an average success rate of 59.6% on iTHOR, clearly above the 31.3% of AVDC and the 34.2% of VideoAgent. The largest improvements appear in Kitchen with 88.3% and in Bedroom with 63.3%, where cluttered and variable layouts require precise reasoning about object positions and feasible motion. Living Room also benefits from stronger spatial grounding, reaching 50.0%. Bathroom performance is similar to the baselines due to its highly constrained geometry, which limits the impact of planning. Overall, the results indicate that explicit spatial structure, rather than relying solely on visual priors, is crucial for generalizing across diverse room configurations.
The qualitative rollouts in Fig. 3 further illustrate this advantage. Baseline trajectories drift because diffusion priors alone cannot preserve stable spatial alignment, leading to early pose errors in long-horizon tasks such as Painting and Mirror. This issue is amplified in iTHOR, where frequent camera rotations require the model to maintain scene consistency across changing viewpoints. Our method preserves the global layout more faithfully: generated frames remain coherent after rotation, and the target is localized consistently throughout the sequence. By encoding spatial plans that constrain depth reasoning, object–agent relations, and admissible motion, the model produces trajectories that better reflect the actual room topology even under viewpoint shifts.
4.3 Real-World Experiments
We evaluate our system in a real-world setting using an SO-100 robotic arm [so100_wiki] on a pick-doll task, where the robot must grasp a yellow rabbit toy and place it into a box. Since real-world data lack depth, we estimate depth with DepthAnything [depthanything] and feed the RGB frame, predicted depth, and object spatial information into the spatial plan table. For action prediction, we use the top-down current and goal frames together with the wrist-camera frame to improve spatial grounding. As shown in Figure 5, we evaluate 20 trials and achieve a 30% success rate, indicating that our model can still complete the task even under incomplete visual observations and depth estimation noise.
Figure 5: Real-world execution of the pick doll task.
4.4 Ablation Study
@@
Ablation on the Video Generation Module. We evaluate three strategies in the video generation module in Meta-World. (1) The Baseline follows AVDC and conditions the diffusion model only on the task prompt and initial frame. (2) Our Distance variant adds the spatial offset between the end effector and the target as raw spatial guidance. (3) Our Subplan variant transforms this offset into structured action cues using a vision language model and injects them as high level spatial guidance As shown in Table 3, both spatial variants outperform the AVDC baseline, demonstrating the utility of explicitly modeling spatial structures for enhanced visuomotor performance. Moreover, the subplan-based approach achieves further gains by abstracting raw spatial cues into structured guidance, highlighting the importance of structured spatial abstraction in facilitating precise and actionable visual plan generation.
Ablation on the Action Prediction Module. We compare two action prediction strategies in Meta-World. (1) AVDC converts optical flow between the current and goal frames into a target coordinate and executes it with a classical controller and heuristic gripper rules. (2) Our method instead uses a flow-based diffusion policy that directly predicts actions from the current frame, goal frame, flow, and robot spatial coordinates, allowing spatial and temporal cues to be modeled jointly. As shown in Table 4, our flow-based diffusion policy consistently outperforms AVDC across all difficulty levels. The average success rate rises from 19.6% for AVDC to 63.4% for our method, with especially large gains on Medium, Hard, and Very Hard tasks. These results highlight that learning based spatial reasoning from flow yields more reliable and adaptable action generation than mapping flow fields to target coordinates through hand-designed controllers.
Ablation on Spatial Condition Injection Strategy in Video Generation. We compare two strategies for injecting spatial plans into the diffusion model. (1) The Global Cond variant concatenates the spatial plan table with the task embedding and applies it uniformly to all U-Net layers. (2)The Local Cond variant follows the GLIGEN formulation and injects the spatial plan only into intermediate layers through cross attention, enabling more localized spatial modulation. As shown in Table 5, Local Conditioning yields higher success rates, reaching 84.3% on average and improving performance on harder tasks such as Hard and Very Hard. This suggests that localized spatial injection improves spatial precision in video synthesis and leads to more reliable downstream control.
Table 3: Comparison of task success rates (%) using different video generation strategies across four levels of task difficulty in Meta-World. AVDC uses text and initial frame only; Distance adds raw spatial offsets; Subplan injects structured spatial subplans.
Easy Medium Hard Very Hard Average
@@
Subplan 94.8% 69.3% 21.3% 78.7% 82.0%
Table 4: Comparison of task success rates (%) across four difficulty levels in Meta-World. AVDC refers to AVDC’s action predictor that directly maps optical flow to actions. Ours uses a flow-based diffusion policy.
Easy Medium Hard Very Hard Average
AVDC 25.9% 10.7% 5.3% 8.0% 19.6%
@@
Table 5: Comparison of task success rates (%) using different spatial condition injection strategies for video generation across four levels of task difficulty in Meta-World. Global Cond injects the spatial plan table across all U-Net blocks; Local Cond follows GLIGEN-style partial injection via cross-attention at intermediate layers.
Easy Medium Hard Very Hard Average
Global Cond 94.8% 63.3% 21.3% 78.7% 82.0%
@@
Ablation on the Replanning Mechanism. We compare three replanning settings: (1) no replanning, (2)replanning only during video generation, and (3) replanning during both video generation and real execution. As shown in Table 6, performance improves as more stages incorporate replanning. Replanning in both imagination and execution achieves the highest average success rate of 86.7%, with clear gains on harder tasks. While generation only replanning offers moderate improvement, combining it with execution time replanning enables the system to correct failures, adapt to environmental changes, and stabilize long horizon behavior. Without replanning, the policy cannot recover from unexpected variations. These results indicate that using replanning in both stages is essential for robust long horizon control.
Table 6: Comparison of task success rates (%) using different replanning mechanisms across four levels of task difficulty in Meta-World. NR denotes no replanning; GR denotes generation replanning; ER denotes execution replanning. Bold indicates best.
Easy Medium Hard Very Hard Average
NR 94.8% 63.3% 21.3% 78.7% 82.0%
@@
Ablation on Finetuning with Expert-Assisted Recovery Data. We evaluate two training configurations. The Original variant uses only fully successful demonstrations, while the Expert-Assisted variant is finetuned on additional trajectories where failures are corrected by an expert policy. As shown in Table 7, expert assisted finetuning improves performance across all difficulty levels, raising the average success rate from 82.0% to 85.6% and yielding substantial gains on Medium and Hard tasks. These results indicate that exposure to partial-to-successful executions enhances robustness by covering recovery scenarios absent in the original dataset. Overall, expert assisted augmentation provides a simple and effective boost to long-horizon reliability.
Table 7: Comparison of task success rates (%) with and without expert assisted data augmentation across four difficulty levels in Meta-World. Original denotes the model trained only on the standard dataset. Expert-Assisted denotes the model additionally trained on expert-corrected rollouts.
Easy Medium Hard Very Hard Average
Original 94.8% 63.3% 21.3% 78.7% 82.0%
Expert-Assisted 96.5% 74.9% 38.7% 84.0% 85.6%
4.5 Error Type Analysis
Human evaluation on MetaWorld reveals that 65% of the generated videos are correct, with errors mainly attributed to fluency issues (9.1%), robot inconsistency (11.0%), and object inconsistency (14.5%). For action execution, 85.5% of the predictions are correct, with the remaining errors caused by physics violations (9.1%) and spatial mistakes (5.5%). The low spatial error rate indicates that the spatial planning module and spatial-conditioned policy effectively capture task geometry, and most failures arise from challenging contact dynamics rather than misinterpreted spatial relations.
4.6 Robustness to Occlusions via Masking
To evaluate the robustness of our video generation module under partial visual occlusions, we introduce a masking mechanism before feeding the current frame into the diffusion model. Specifically, we randomly mask a portion of the image input, simulating occlusions such as those caused by environmental factors or sensor noise.
4.7 Hyperparameter Analysis
@@
Effect of Hyperparameter in Action Prediction. We analyze key design choices in the action prediction module. First, we study the effect of goal frame selection range. Setting the goal frame as a random future frame within the next $k=20$ timesteps yields over 10% higher average success rates compared to restricting it to a narrower window between 17–20 timesteps ahead, indicating that allowing a broader temporal horizon improves generalization and task adaptability. Second, regressing the noise $\epsilon$ achieves better performance than directly predicting the denoised action $\hat{a}_{0}$, which contrasts with some prior findings [diffusionpolicy]. Lastly, while direct MLP-based action regression converges faster, the diffusion-based denoising model demonstrates superior performance and robustness across tasks.
5 Conclusion
We present SP, a spatial-aware visuomotor framework that integrates explicit spatial modeling and reasoning into robot manipulation. Central to our approach is a spatial plan table that encodes atomic actions, directions, and distances, enabling structured spatial guidance throughout the pipeline. Our policy generates spatially coherent visual trajectories by conditioning a video diffusion model on a spatial plan table, which are then used to guide action prediction based on temporally aligned spatial coordinates. A dual-stage replanning mechanism enhances spatial reasoning by refining the spatial plan table and regenerating video and actions upon inconsistency, ensuring closed-loop correction. Experiments on MetaWorld show that SP improves success rates by over 33% compared to prior methods, validating the effectiveness of spatial modeling and reasoning.
[…]
Supplementary Material
A.1 Inference Pipeline